# LocDiffusion: Identifying Locations on Earth by Diffusing in the Hilbert Space

### v1

@@ A.1 Sparsity of Existing Positional Encoding Methods
A.2 Computation of Spherical Harmonics
@@
A.3 Spatial Resolution of SHDD Encoding/Decoding
The spatial resolution of SHDD encoding/decoding (i.e., on what scales the mode-seeking decoder can accurately locate the probability mass concentration of the spherical Dirac delta functions) is bound by the degree $L$ of Legendre polynomials. For an $L$-degree SHDD representation, the spatial scale threshold at which it can accurately approximate spherical functions is $\pi/L$ in radian or approximately $20000/L$ in kilometers (Ince et al., 2019). For example, for $L=15,23,31$, the thresholds are around 1300 km, 870 km, and 640 km, respectively. At scales significantly below half this threshold, even if the diffusion model generates accurate coefficient vectors, the mode-seeking decoder can still only decode vague locations with large variances. Figure 3 gives a visual intuition.
@@ A.3 Spatial Resolution of SHDD Encoding/Decoding
Therefore, to uplift the performance of LocDiffusion, one straightforward way is to use larger $L$. We conduct an ablation study of the effect of $L$ on image geolocalization performance on the Im2GPS3K dataset. The results are shown in Figure 4. We can see as $L$ increases, while the model performances at larger spatial scales (e.g., 750km, 2500km) only increase slightly, the performances at smaller scales (e.g., 1km, 25km, 200km) see huge uplifting. This validates our hypothesis – a larger $L$ can make the mode-seeking decoder decoding vague locations with smaller variances, thus leading to higher image geolocalization performance. The largest $L$ we tried in Figure 4 is 47 which corresponds to a spatial resolution of $200km$. This is why we see huge performance improvements on the $25km$ and $200km$ curves but not on the $1km$ curve since $1km$ is still significantly smaller than the current spatial scale threshold.
However, it is not recommended to unlimitedly increase $L$. There are two major reasons:
  1. 1.
@@
    The SHDD encoding dimension increases quadratically with $L$, i.e., we need quadratic space to halve the spatial resolution. It is expensive and difficult to train a diffusion model on very large encodings (e.g. to achieve 50 km spatial resolution, we theoretically need 160,000 dimensions). We find that the higher the dimension of SHDD encodings, the higher the maximum absolute values of the coefficients. Figure 4 is a log-scale plot of the maximum absolute values of each SHDD encoding dimension up to $L=63$ (i.e., in total 64 $\times$ 64 = 4096 dimensions). The absolute values below 2500 dimensions are in general manageable with only a few spikes. However, dimensions beyond this threshold become unbearably large, which makes the probability computation very unstable and easy to overflow.
Based on these observations, we use up to $L=47$ in our paper because now the dimension of SHDD encoding goes to $2304$, still within the manageable range.
@@
Figure 4: (a): An illustration of how the image geolocalization performance on the Im2GPS3K dataset increases as L increases. Different curves indicate performance metrics on different spatial scales. (b): A log-scale plot of the maximum absolute values of each SHDD encoding dimension up to 64$\times$64 = 4096 dimensions.
@@
Moreover, to address the high dimension issue when we use a large $L$, we find that applying a low-pass filter to the dimensions is a good dimension reduction solution. See Figure 4. Many dimensions of the SHDD encodings have very small absolute values and will not significantly influence the results of SHDD encoding/decoding. Thus, we can set a low-pass filter analogous to Fourier transformation and signal processing, which only keeps the dimensions that have adequately large coefficient values.
A.4 A LocDiffusion+GeoCLIP Hybrid Appoach
@@
From Table 1, we can see that LocDiffusion is a generative model and outperforms GeoCLIP on coarser scales (¿ 200 km) while GeoCLIP outperforms on finer scales (¡25km). Inspired by this, we develop a hybrid approach which uses LocDiffusion’s predictions to narrow down the candidate regions and then deploy GeoCLIP to generate the final predictions. More specifically, we first use LocDiffusion to sample multiple times (e.g., 16) and get a rough distribution of candidate locations, i.e. they indicate where the true answer is highly likely to reside. Then, we restrict the retrieval of GeoCLIP to the neighborhoods (200 km radius) of these candidate locations. We find that this hybrid approach yields substantiate better results than both GeoCLIP and LocDiffusion alone. This approach is similar to recommendation systems’ retrieve and rerank approach. This flexible hybrid approach points to an interesting future research direction. We can also replace GeoCLIP with other state-of-the-art image geolocalization models such as PIGEON to further improve the model performance.
A.5 Inductive Bias of Gallery
@@
Figure 5 is a set of visualizations of Table 2. It clearly demonstrates how GeoCLIP suffers greatly from using a grid gallery without prior knowledge (i.e., without using the inductive bias brought by the MP16 image gallery), while our method remains almost unaffected on larger spatial scales (200 km, 750 km, and 2500 km) and much less affected on smaller scales (25 km). These results clearly demonstrate that the high performance of GeoCLIP on smaller spatial scales is based on the fact that the MP16 image gallery used by GeoCLIP already contains candidate locations that are close enough to true answers (i.e., test image locations). However, this is not the case for our method because our model does not rely on such an image gallery either during training or during inferencing time. Thus, our LocDiffusion model suffers much less when we switch to a grid location gallery. Moreover, when we decrease the number of points in the grid gallery, the performances of GeoCLIP decrease significantly while the performances of our LocDiffusion are almost unaffected.
@@ A.6 Computational Complexity
The major factor that decides the inference time of LocDiffusion is the choice of the sampler. In our experiment, we use the original DDPM sampler (i.e., no DDIM acceleration) with 100 sampling steps. The inference time per image for LocDiffusion is 0.056s and for GeoCLIP 0.024 seconds.
A.7 Ablation Studies
A.7.1 Comparison with other location encoding/decoding techniques
As we have discussed, the superiority of using SHDD for location decoding is that its encoding space is smoother than other location encoders that use neural networks such as rbf and Sphere2Vec (Mai et al., 2023b). To demonstrate this, we evenly sample 1 million locations on Earth, encode them into corresponding location embeddings by using rbf and Sphere2Vec location encoder, and train a neural network decoder to map the location embeddings back to locations. We also use the learned neural decoder in the LocDiffusion training with weights frozen. The ablation study results are shown in Table 5. We can see that the performances of rbf and Sphere2Vec are much worse than SHDD, especially on smaller scales. This is because: (1) the learned decoder is not 100% accurate, i.e. it may decode an encoding to a wrong location, and (2) if the encoding gets a small perturbation, the decoded location may have a very large drift due to non-linearity.
@@ A.7.1 Comparison with other location encoding/decoding techniques
rbf (Mai et al., 2020a) NN 0.0 0.0 18.2 44.1 60.2
Sphere2Vec (Mai et al., 2023b) NN 0.0 0.0 22.1 58.4 72.3
SHDD SHDD 10.9 34.0 53.3 72.5 85.2
@@
SHDD SHDD 5.3 km
A.7.2 Comparison with other losses
@@
A.7.3 Ablation studies on other modules
@@
A.8 Training Hyper parameter Setup

### v2

@@ A.2 Sparsity of Existing Positional Encoding Methods
A.3 Computation of Spherical Harmonics
@@
A.4 Spatial Resolution of SHDD Encoding/Decoding
The spatial resolution of SHDD encoding/decoding (i.e., on what scales the mode-seeking decoder can accurately locate the probability mass concentration of the spherical Dirac delta functions) is bound by the degree $L$ of Legendre polynomials. For an $L$-degree SHDD representation, the spatial scale threshold at which it can accurately approximate spherical functions is $\pi/L$ in radian or approximately $20000/L$ in kilometers [16]. For example, for $L=15,23,31$, the thresholds are around 1300 km, 870 km, and 640 km, respectively. At scales significantly below half this threshold, even if the diffusion model generates accurate coefficient vectors, the mode-seeking decoder can still only decode vague locations with large variances. Figure 4 gives a visual intuition.
@@ A.4 Spatial Resolution of SHDD Encoding/Decoding
Therefore, to uplift the performance of LocDiff, one straightforward way is to use larger $L$. We conduct an ablation study of the effect of $L$ on image geolocalization performance on the Im2GPS3K dataset. The results are shown in Figure 5. We can see as $L$ increases, while the model performances at larger spatial scales (e.g., 750km, 2500km) only increase slightly, the performances at smaller scales (e.g., 1km, 25km, 200km) see huge uplifting. This validates our hypothesis – a larger $L$ can make the mode-seeking decoder decoding vague locations with smaller variances, thus leading to higher image geolocalization performance. The largest $L$ we tried in Figure 5 is 47 which corresponds to a spatial resolution of $200km$. This is why we see huge performance improvements on the $25km$ and $200km$ curves but not on the $1km$ curve since $1km$ is still significantly smaller than the current spatial scale threshold.
@@
However, it is not recommended to unlimitedly increase $L$. There are two major reasons: The SHDD encoding dimension increases quadratically with $L$, i.e., we need quadratic space to halve the spatial resolution. It is expensive and difficult to train a diffusion model on very large encodings (e.g. to achieve 50 km spatial resolution, we theoretically need 160,000 dimensions). We find that the higher the dimension of SHDD encodings, the higher the maximum absolute values of the coefficients. Figure 5 is a log-scale plot of the maximum absolute values of each SHDD encoding dimension up to $L=63$ (i.e., in total 64 $\times$ 64 = 4096 dimensions). The absolute values below 2500 dimensions are in general manageable with only a few spikes. However, dimensions beyond this threshold become unbearably large, which makes the probability computation very unstable and easy to overflow. Based on these observations, we use up to $L=47$ in our paper because now the dimension of SHDD encoding goes to $2304$, still within the manageable range.
@@
Figure 5: (a): An illustration of how the image geolocalization performance on the Im2GPS3K dataset increases as L increases. Different curves indicate performance metrics on different spatial scales. (b): A log-scale plot of the maximum absolute values of each SHDD encoding dimension up to $\times64$ = 4096 dimensions.
@@
Moreover, to address the high dimension issue when we use a large $L$, we find that applying a low-pass filter to the dimensions is a good dimension reduction solution. See Figure 5. Many dimensions of the SHDD encodings have very small absolute values and will not significantly influence the results of SHDD encoding/decoding. Thus, we may set a low-pass filter analogous to Fourier transformation and signal processing, which only keeps the dimensions that have adequately large coefficient values. This can be further investigated in future works.
A.5 Pseudo-Code for LocDiff
Algorithm 1 Training LocDiff
Input :
A dataset $\mathcal{D}$ with location-image pairs $\{(\mathbf{p},\mathbf{I})\}$. Each location is a tuple of latitude and longitude $\mathbf{p}=(\theta,\phi)$. A Gaussian noise scheduler $\mathcal{N}(t)$, where $t$ is the time step. SHDD encoder $\mathbb{PE}_{\text{SHDD}}$. Pretrained Image encoder $\mathbb{E}_{\text{Im}}$. CS-UNet model $M$ with random initialization. SHDD KL-divergence loss $\mathcal{L}_{\text{SHDD-KL}}$.
Output :
A trained CS-UNet model $M$.
For $\mathbf{p},\mathbf{I}\in\mathcal{D}$:
compute the SHDD encoding of $\mathbf{p}$: $\mathbf{e}\leftarrow\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$;
compute the image embedding of $\mathbf{I}$: $\mathbf{e_{I}}\leftarrow\mathbb{E}_{\text{Im}}(\mathbf{I})$;
randomly draw a time step $t$;
add Gaussian noise to the SHDD encoding (forward process): $\mathbf{e}^{\prime}\leftarrow\mathbf{e}+\mathcal{N}(t)$;
use the CS-UNet to denoise the SHDD encoding conditioned on the image embedding (backward process): $\hat{\mathbf{e}}\leftarrow M(\mathbf{e}^{\prime},\mathbf{e_{I}},t)$;
compute the SHDD KL-divergence loss: $l\leftarrow\mathcal{L}_{\text{SHDD-KL}}(\hat{\mathbf{e}},\mathbf{e})$;
use gradient decent to minimize $l$ and update $M$: $M\leftarrow\operatorname*{arg\,min}_{M}l$;
return $M$
Algorithm 2 Inferencing LocDiff
Input :
Image $\mathbf{I}$. Random image augmentation $\mathbb{AUG}$. Ensemble number $N$. DDPM sampler $\mathbb{DDPM}$. DDPM step $T$. Pretrained Image encoder $\mathbb{E}_{\text{Im}}$. Trained CS-UNet model $M$ with random initialization. SHDD mode-seeking decoder $\mathbb{PD}_{\text{mode}}$. Mode-seeking range hyperparameter $\rho$. An initial location $\mathbf{p}=(0,0)$.
Output :
An ensembled location prediction $\hat{\mathbf{p}}$.
For $i=1,2,\cdots N$:
generate a random augmentation of image $\mathbf{I}$: $\mathbf{I}_{\text{aug}}\leftarrow\mathbb{AUG}(\mathbf{I})$;
compute the image embedding of $\mathbf{I}$: $\mathbf{e_{\text{aug}}}\leftarrow\mathbb{E}_{\text{Im}}(\mathbf{I}_{\text{aug}})$;
randomly draw an SHDD encoding $\mathbf{e}^{\prime}$;
use the DDPM sampler and the trained CS-UNet $M$ to sample an SHDD encoding conditioned on the image embedding: $\hat{\mathbf{e}}\leftarrow\mathbb{DDPM}(M,\mathbf{e}^{\prime},\mathbf{e_{\text{aug}}},T)$;
use the SHDD decoder $\mathbb{PD}_{\text{mode}}$ to decode a location from the sampled SHDD encoding: $\hat{\mathbf{p}}\leftarrow\mathbb{PD}_{\text{mode}}(\hat{\mathbf{e}},\rho)$;
$\mathbf{p}\leftarrow\mathbf{p}+\hat{\mathbf{p}}$
return the ensemble of $N$ location predictions: $\mathbf{p}/N$
A.6 A LocDiff-H Hybrid Appoach
@@
We develop a hybrid approach which uses LocDiff’s predictions to narrow down the candidate regions and then deploy GeoCLIP to generate the final predictions. More specifically, we first use LocDiff to sample multiple times (e.g., 16) and get a rough distribution of candidate locations, i.e. they indicate where the true answer is highly likely to reside. Then, we restrict the retrieval of GeoCLIP to the neighborhoods (200 km radius) of these candidate locations. As shown in Table 1, we can see that this hybrid approach yields substantially better results than both GeoCLIP and LocDiff alone. This approach is similar to the recommendation systems’ retrieve and rerank approach. This flexible hybrid approach points to an interesting future research direction. We can also replace GeoCLIP with other state-of-the-art image geolocalization models such as PIGEON [10], to further improve the model performance.
A.7 Inductive Bias of Gallery
@@
Figure 6 is a set of visualizations of Table 3. It clearly demonstrates how GeoCLIP suffers greatly from using a grid gallery without prior knowledge (i.e., without using the inductive bias brought by the MP16 image gallery), while our method remains almost unaffected on larger spatial scales (200 km, 750 km, and 2500 km) and much less affected on smaller scales (25 km). These results clearly demonstrate that the high performance of GeoCLIP on smaller spatial scales is based on the fact that the MP16 image gallery used by GeoCLIP already contains candidate locations that are close enough to true answers (i.e., test image locations). However, this is not the case for our method because our model does not rely on such an image gallery either during training or during inferencing time. Thus, our LocDiff model suffers much less when we switch to a grid location gallery. Moreover, when we decrease the number of points in the grid gallery, the performances of GeoCLIP decrease significantly while the performances of our LocDiff are almost unaffected.
@@ A.8 Computational Complexity
The major factor that decides the inference time of LocDiff is the choice of the sampler. In our experiment, we use the original DDPM sampler (i.e., no DDIM acceleration) with 200 sampling steps. The inference time per image for LocDiff is 0.056s and for GeoCLIP 0.024 seconds.
A.9 Ablation Studies
A.9.1 Comparison with other location encoding/decoding techniques
As we have discussed, the superiority of using SHDD for location decoding is that its encoding space is smoother than other location encoders that use neural networks such as rbf and Sphere2Vec [27]. To demonstrate this, we evenly sample 1 million locations on Earth, encode them into corresponding location embeddings by using rbf and Sphere2Vec location encoder, and train a neural network decoder to map the location embeddings back to locations. We also use the learned neural decoder in the LocDiff training with weights frozen. The ablation study results are shown in Table 7. We can see that the performances of rbf and Sphere2Vec are much worse than SHDD, especially on smaller scales. This is because: (1) the learned decoder is not 100% accurate, i.e. it may decode an encoding to a wrong location, and (2) if the encoding gets a small perturbation, the decoded location may have a very large drift due to non-linearity.
@@ A.9.1 Comparison with other location encoding/decoding techniques
rbf [24] NN 0.0 0.0 18.2 44.1 60.2
Sphere2Vec [27] NN 0.0 0.0 22.1 58.4 72.3
SHDD (L=47) SHDD (L=47) 10.9 34.0 53.3 72.5 85.2
@@
SHDD (L=47) SHDD (L=47) 5.3 km
A.9.2 Comparison with other losses
@@
A.9.3 Ablation studies on other modules
@@
A.10 Training Hyper parameter Setup